OpenStudy (anonymous):

**Medal to everyone who gives a well explained answer Directions: Use the pythagorean theorem to find th distance, to the nearest tenth, between each pair of points. 11. *Shows a graph* *On the graph Y is on (0,1) and Z is on (4.5).

OpenStudy (anonymous):

 |AB|=(4,5)-(0,1)=(4,4) according to brother pathagoris thats nothing more than the SQRT(4^2+4^2)=SQRT(32)=4SQRT(2)